X + 2Y + Z = 22
X + Y = 15
3X + Y + Z = -37

Answer by user_dude2008(1862)

X + 2Y + Z = 22
X + Y = 15 --------> Y=15-X
3X + Y + Z = -37



X + 2Y + Z = 22
X + 2(15-X) + Z = 22
X+30-2X+Z=22
-X+Z=-8
Z=X-8


3X + Y + Z = -37
3X + 15-X + X-8 = -37
3X+7=-37
3X=-44
X=-44/3

Y=15-X=15-(-44/3)=89/3
Y=89/3

Z=X-8=-44/3-8=-68/3
Z=-68/3